# Edit3K: Universal Representation Learning for Video Editing Components

### v1

@@ Abstract
This paper focuses on understanding the predominant video creation pipeline, i.e., compositional video editing with six main types of editing components, including video effects, animation, transition, filter, sticker, and text. In contrast to existing visual representation learning of visual materials (i.e., images/videos), we aim to learn visual representations of editing actions/components that are generally applied on raw materials. We start by proposing the first large-scale dataset for editing components of video creation, which covers about $3,094$ editing components with $618,800$ videos. Each video in our dataset is rendered by various image/video materials with a single editing component, which supports atomic visual understanding of different editing components. It can also benefit several downstream tasks, e.g., editing component recommendation, editing component recognition/retrieval, etc. Existing visual representation methods perform poorly because it is difficult to disentangle the visual appearance of editing components from raw materials. To that end, we benchmark popular alternative solutions and propose a novel method that learns to attend to the appearance of editing components regardless of raw materials. Our method achieves favorable results on editing component retrieval/recognition compared to the alternative solutions. A user study is also conducted to show that our representations cluster visually similar editing components better than other alternatives. Furthermore, our learned representations used to transition recommendation tasks achieve state-of-the-art results on the AutoTransition dataset. The code and dataset will be released for academic use.
footnotetext: *** Corresponding author, sijiezhu@bytedance.comfootnotetext: $\dagger$ Equal contributionfootnotetext: $\ddagger$ This work was done during the first author’s internship at ByteDance
@@ 2 Related Work
[…]
@@ 5 Experiment
[…]
@@ 5.3 Editing Components Representation Distribution
[…]
@@
The users are given 5 videos in a row, where the left column is the video for the query editing component, and the other columns show the example videos for the retrieved editing components corresponding to the $4$ methods (an example is shown in Fig. 5). The users are asked to select the videos that are visually similar to the query and each query is evaluated by at least $10$ users. The selected results are considered satisfactory and we compute the satisfactory rate of different methods on all queries. As shown in Tab. 5.3, ‘Ours’ significantly outperforms the other three methods in terms of average satisfactory rate, indicating the effectiveness of the proposed guidance architecture and additional loss on embedding distribution. The result is also consistent across different editing component categories. Qualitative results is provided in Fig. 6. The t-SNE [30] visualization for the embedding space is provided in the supplementary material.
[…]
@@
[…]
@@ 5.4 Ablation Study
Different Modules. We conduct ablation study in Tab. 5.3 by removing different modules to demonstrate the effectiveness of each module. The ‘Baseline’ model with only spatial-temporal encoder and in-batch InfoNCE loss achieves an average R@1 score of $31.2\%$. Adding the embedding queue loss term (‘Baseline + $\mathcal{L}_{q}$’) brings $5.4\%$ improvement on the R@1 accuracy, indicating the effectiveness of the $\mathcal{L}_{q}$ in Eq. 2. The R@1 is further improved by $3.4\%$ when adding guidance tokens in the spatial encoder (‘Baseline + $\mathcal{L}_{q}$ + GT’). Instead, deploying the guided embedding decoder (‘Baseline + $\mathcal{L}_{q}$ + GD’) also gains $12.4\%$ improvement on R@1 over ‘Baseline + $\mathcal{L}_{q}$’. The performance is further improved when all the modules are employed (‘Baseline + $\mathcal{L}_{q}$ + GT + GD’), demonstrating the effectiveness of the guidance architecture.
@@ 5.5 Transition Recommendation
We conduct experiments on transition recommendation as a downstream task to validate the generalization ability of our learned universal representation. In [27], the embeddings of transitions are pre-computed with a classification model trained on AutoTransition dataset, and the embeddings are fixed during the training of the recommendation model. As shown in Fig. 5.3, we follow [27] to train transition recommendation models which take a raw video as input and generate ranking scores for a set of transitions. We replace the pre-computed transition embeddings with other alternative methods to validate the benefit of a strong representation on downstream tasks. We follow [27] to report R@k and mean rank for evaluation.
@@
Tab. 5.3 shows that the embedding of the proposed method trained on AutoTransition datasets (‘Ours-A’) outperforms the state-of-the-art result in [27] (‘Classification’), indicating the superiority of the proposed architecture. Since the transitions in AutoTransition dataset are included in our Edit3K dataset, we further adopt the pre-computed embedding on Edit3K (‘Ours-E’) and it brings additional performance boosts, indicating the superiority of our diverse and balanced dataset. Note that all the rows use the same recommendation method and the only difference is the pre-computed transition embeddings, a slight performance boost could indicate significant improvement on representation learning.

### v2

@@ Abstract
This paper focuses on understanding the predominant video creation pipeline, i.e., compositional video editing with six main types of editing components, including video effects, animation, transition, filter, sticker, and text. In contrast to existing visual representation learning of visual materials (i.e., images/videos), we aim to learn visual representations of editing actions/components that are generally applied on raw materials. We start by proposing the first large-scale dataset for editing components of video creation, which covers about $3,094$ editing components with $618,800$ videos. Each video in our dataset is rendered by various image/video materials with a single editing component, which supports atomic visual understanding of different editing components. It can also benefit several downstream tasks, e.g., editing component recommendation, editing component recognition/retrieval, etc. Existing visual representation methods perform poorly because it is difficult to disentangle the visual appearance of editing components from raw materials. To that end, we benchmark popular alternative solutions and propose a novel method that learns to attend to the appearance of editing components regardless of raw materials. Our method achieves favorable results on editing component retrieval/recognition compared to the alternative solutions. A user study is also conducted to show that our representations cluster visually similar editing components better than other alternatives. Furthermore, our learned representations used to transition recommendation tasks achieve state-of-the-art results on the AutoTransition dataset. The code and dataset are released at https://github.com/GX77/Edit3K.
footnotetext: * Corresponding author, sijiezhu@bytedance.comfootnotetext: $\dagger$ Equal contributionfootnotetext: $\ddagger$ This work was done during the first author’s internship at ByteDance
@@ 2 Related Work
Video Effect Animation Transition Filter Sticker Text Total
@@ 5 Experiment
Video Effect Animation Transition Filter Sticker Texts Avg.
R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10
@@ 5.3 Editing Components Representation Distribution
Method V. E. Anim. Tran. Filt. Stic. Text Avg.
@@
The users are given 5 videos in a row, where the left column is the video for the query editing component, and the other columns show the example videos for the retrieved editing components corresponding to the $4$ methods (an example is shown in Fig. 5). The users are asked to select the videos that are visually similar to the query and each query is evaluated by at least $10$ users. The selected results are considered satisfactory and we compute the satisfactory rate of different methods on all queries. As shown in Tab. 3, ‘Ours’ significantly outperforms the other three methods in terms of average satisfactory rate, indicating the effectiveness of the proposed guidance architecture and additional loss on embedding distribution. The result is also consistent across different editing component categories. Qualitative results is provided in Fig. 6. The t-SNE [30] visualization for the embedding space is provided in the supplementary material.
Video Effect Animation Transition Filter Sticker Texts Avg.
R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10
@@
Method R@1 R@10
None 36.6 73.8
K-Means 52.1 79.9
Default 6 Types 52.6 80.6
Table 5: Ablation of different methods for generating embedding centers on Edit3K for editing components retrieval in terms of R@k (%).
@@
Transition Embedding Performance
Method Dataset R@1 R@5 Rank ($\downarrow$)
@@ 5.4 Ablation Study
Different Modules. We conduct ablation study in Tab. 4 by removing different modules to demonstrate the effectiveness of each module. The ‘Baseline’ model with only spatial-temporal encoder and in-batch InfoNCE loss achieves an average R@1 score of $31.2\%$. Adding the embedding queue loss term (‘Baseline + $\mathcal{L}_{q}$’) brings $5.4\%$ improvement on the R@1 accuracy, indicating the effectiveness of the $\mathcal{L}_{q}$ in Eq. 2. The R@1 is further improved by $3.4\%$ when adding guidance tokens in the spatial encoder (‘Baseline + $\mathcal{L}_{q}$ + GT’). Instead, deploying the guided embedding decoder (‘Baseline + $\mathcal{L}_{q}$ + GD’) also gains $12.4\%$ improvement on R@1 over ‘Baseline + $\mathcal{L}_{q}$’. The performance is further improved when all the modules are employed (‘Baseline + $\mathcal{L}_{q}$ + GT + GD’), demonstrating the effectiveness of the guidance architecture.
@@ 5.5 Transition Recommendation
We conduct experiments on transition recommendation as a downstream task to validate the generalization ability of our learned universal representation. In [27], the embeddings of transitions are pre-computed with a classification model trained on AutoTransition dataset, and the embeddings are fixed during the training of the recommendation model. As shown in Fig. 7, we follow [27] to train transition recommendation models which take a raw video as input and generate ranking scores for a set of transitions. We replace the pre-computed transition embeddings with other alternative methods to validate the benefit of a strong representation on downstream tasks. We follow [27] to report R@k and mean rank for evaluation.
@@
Tab. 6 shows that the embedding of the proposed method trained on AutoTransition datasets (‘Ours-A’) outperforms the state-of-the-art result in [27] (‘Classification’), indicating the superiority of the proposed architecture. Since the transitions in AutoTransition dataset are included in our Edit3K dataset, we further adopt the pre-computed embedding on Edit3K (‘Ours-E’) and it brings additional performance boosts, indicating the superiority of our diverse and balanced dataset. Note that all the rows use the same recommendation method and the only difference is the pre-computed transition embeddings, a slight performance boost could indicate significant improvement on representation learning.